2a+3b=7
5a-4b=6



Answer by tommyt3rd(5050)

m4: 2a+3b=7
m3: 5a-4b=6
8a+12a=28
15a-12a=18
23a=46
a=2

2(2)+3b=7
3b=3
b=1

:)